Opened 9 years ago

Closed 9 years ago

Last modified 3 years ago

#342 closed flight processing (fixed)

GB08/06, flight day 167/2010, Bleaklow

Reported by: ella Owned by: knpa
Priority: alpha 4 low Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: iopa@… Other processors: mark1, benj

Description (last modified by emca)

Location of unpacked data: /users/rsg/arsf/workspace/GB08_06-2010_167_Bleaklow

Scientific purpose: Study of peatlands in the uplands of the Peak District National Park.

Scientific objectives: To evaluate the ability of SPECIM images to distinguish between moorland restoration treatments and provide new
knowledge on thresholds for estimating habitat condition for sparse vegetation growth on carbon-rich soils, building on work on hyperspectral remote sensing of bare peat.

Data arrived from ARSF via network transfer on 28th June 2010.

PI: Ms Julia McMorrow

Sensors:

  • Eagle (18/08/10)
  • Hawk (20/12/10)
  • Leica LIDAR (18/08/10)
  • RCD Digital photography (24/08/10)

Change History (52)

comment:1 Changed 9 years ago by ella

  • Component changed from ARSF to Processing: general

comment:2 Changed 9 years ago by benj

  • Summary changed from BG08/06, flight day 167/2010, Bleaklow to GB08/06, flight day 167/2010, Bleaklow

comment:3 Changed 9 years ago by benj

  • Description modified (diff)

comment:4 Changed 9 years ago by ella

This project has not yet been fast copied into the arsf repoistry.

comment:5 Changed 9 years ago by ella

Applanix navigation processing done.

comment:6 Changed 9 years ago by ella

IPAS navigation processing is done.

comment:7 Changed 9 years ago by jb362

  • Owner changed from mggr to jb362
  • Status changed from new to assigned

Starting hyperspectral processing.

comment:8 Changed 9 years ago by jb362

North end of eagle flightline 8 will not line up with other flightlines.The 0.03 offset seems to be the best fit for the rest of the flightline.

comment:9 Changed 9 years ago by jb362

Playing around with the gpt offsets doesn't appear to sort out the above problem

comment:10 Changed 9 years ago by jb362

I think I may just cut off the top of flight line 8 off since changing the sct or gpt offests doesn't have any effect

comment:11 Changed 9 years ago by benj

No, you shouldn't ever cut data off the flightline. If it can't be rectified completely then put a comment to that effect in the readme prior to delivery, but the data itself is still fine even if not quite in the right place.

comment:12 Changed 9 years ago by jb362

OK will do.

SCT Offsets:

e1: 0.03
e2: 0.01
e3: 0.04
e4: 0.01
e5: 0.00
e6: 0.01
e7: -0.01
e8: 0.03
e9: 0.03
e10: 0.08
e11: 0.05
e12: -0.01
e13: 0.05
e14: 0.02
e15: 0.03
e16: 0.85
e17: 0.01
e18: -0.01
e19: 0.03
e20: 0.04
e21: 0.02
e22: 0.02
e23: 0.05
e24: 0.04
e25: -0.01
e26: 0.03

Sync error on flightline 16 accounts for very large offset

comment:13 Changed 9 years ago by mark1

OS vectors have arrived and have been unpacked to ~arsf/vectors/GB08_06-second

comment:14 Changed 9 years ago by jb362

Flight line e23 split into 4 parts:

e24a:0.05
e24b:0.1
e24c:0.1
e24d:0.06

Flight lines e11, e21, e22 had end of flight line cut off as these sections were out the target area and had large bends resulting in turn errors.

comment:15 Changed 9 years ago by jb362

Typo!

Flight line e23 split into 4 parts:

e23a:0.05
e23b:0.1
e23c:0.1
e23d:0.06

comment:16 Changed 9 years ago by iopa

  • Cc iopa@… added

Taking up LIDAR processing.

comment:17 Changed 9 years ago by jb362

Hawk SCT offsets:

h1: 0.02
h2: 0.01
h3: 0.06
h4: 0.00
h5: 0.03
h6: 1.70
h7: 0.01
h8: 0.02
h9: 0.00
h10: 0.01
h11: 0.02
h12: -0.01
h13: 0.02
h14: 0.03
h15: 0.02
h16: 1.35
h17: 0.00
h18: 0.03
h19: 0.01
h20: 0.03
h21: 0.02
h22: 0.03
h23: 0.02
h24: 0.05
h25: 0.06
h26: 0.01

Sync errors on flight lines 16 & 6 account for very large offsets

comment:18 Changed 9 years ago by iopa

Missing .sup file from IPAS output, will re-process IPAS before proceeding to LIDAR.

comment:19 Changed 9 years ago by jb362

Hyperspectral processing complete and ready for delivery check

comment:20 Changed 9 years ago by anch

Delivery checking the hyperspectral data for this project now.

comment:21 Changed 9 years ago by jb362

Delivery check fixes complete

comment:22 Changed 9 years ago by jb362

Delivery checking LiDAR

comment:23 Changed 9 years ago by jb362

Check complete. Change "Ordance survey" to "Ordnance survey" in data quality remarks.

comment:24 Changed 9 years ago by iopa

Dispatched USB disk containing Eagle and LiDAR on the 19/08/2010.
PI and collaborators have been emailed.

Note: Collaborator 2 will be unavailable and would like any contact regarding remote sensing to be diverted to Stuart Masheder.

Hawk data will be delivered individually when calibration issue has been resolved.

comment:26 Changed 9 years ago by iopa

  • Description modified (diff)

Processing camera.

comment:27 Changed 9 years ago by iopa

Following a request by the PI, first collaborator Aletta Bonn no longer works on the project.
In her stead, Jonathan Walker will act as first collaborator .

Also by request of the PI, Elizabeth Lowe should be CC'ed during any contact.(emails can be found in Internal Contact details page)

comment:29 Changed 9 years ago by jb362

Camera delivery check complete- everything looks good.

comment:30 Changed 9 years ago by iopa

  • Description modified (diff)

comment:31 Changed 9 years ago by benj

  • Other processors set to mark1, benj

Hawk delivery check complete, looks good, ready for delivery.

comment:32 Changed 9 years ago by benj

  • Description modified (diff)

Hawk rsynced back to repository, delivery in post to PI

comment:33 Changed 9 years ago by benj

  • Component changed from Processing: general to Archiving
  • Owner changed from jb362 to knpa

comment:34 Changed 9 years ago by anch

Re-checking hawk level 1 files

comment:35 Changed 9 years ago by anch

Hawk re-checked and ready to deliver

comment:36 Changed 9 years ago by emca

  • Description modified (diff)

Hawk data re-delivered 20/12/2010

comment:37 Changed 9 years ago by emca

Re-processed hawk data has been r-synced, project is ready for archiving.

comment:38 Changed 9 years ago by knpa

Beginning archiving.

comment:39 Changed 9 years ago by knpa

There is some full waveform lidar data in this project.

comment:40 Changed 9 years ago by knpa

There is tabi data for this project.

comment:41 Changed 9 years ago by knpa

This is webcam data for this project.

comment:42 Changed 9 years ago by knpa

Created Tarball and made available to NEODC.

comment:43 Changed 9 years ago by knpa

Archived and backed up by NEODC.
Moved to archived disk.

comment:44 Changed 9 years ago by knpa

  • Milestone set to 2010 data processing completion
  • Resolution set to fixed
  • Status changed from assigned to closed

Deleting workspace version.

Closing ticket.

comment:45 Changed 8 years ago by adbe

Reprocessed eagle level1 files on 30/11/2011, corrected timing offset.

comment:46 Changed 8 years ago by knpa

Reprocessing HS with APL.

From the above it looks like adbe has done the Eagle already.

comment:47 Changed 8 years ago by knpa

Correction - adbe redid the eagle with az in order to fix a problem with the original delivery, apl version has not yet been done.

comment:48 Changed 8 years ago by knpa

Stage 1 completed, running stage 2 for eagle.

comment:49 Changed 8 years ago by knpa

Re-running stage 2 as apl chain has been updated to generate additional masked files.

comment:50 Changed 8 years ago by knpa

Re-ran stage 2 for Hawk as aplcal has been updated in relation to the bad pixel file.

comment:51 Changed 7 years ago by knpa

Stage 3 complete - ready for delivery check.

comment:52 Changed 7 years ago by knpa

Delivery checked by Jan.

Made the recommended changes:

  • mosaics moved to correct place in ReadMe
  • Changed projcode in ReadMe title to GB08/06
  • Changed DEM name to correct format

Putting on disk for Julia.

comment:53 Changed 7 years ago by knpa

delivered to Julia McMorrow in Manchester (25/07/12)

comment:54 Changed 3 years ago by dac

Archiving

Data available from NEODC: http://browse.ceda.ac.uk/browse/neodc/arsf/2010/GB08_06/GB08_06-2010_167_Bleaklow

Tidied up project.

Note: See TracTickets for help on using tickets.